An urn contains 8 red balls. 3 yellow, 9 white, 6 balls

are randomly selected with replacement

What is the probability 2 are red l is yellow, and 3 are

white?
1
Expert's answer
2020-12-24T16:01:06-0500

We can take "2" of "8" red balls by "\binom{8}{2}" ways, "1" of "3" yellow balls by "\binom{3}{1}" ways, and "3" of "9" white balls by "\binom{9}{3}" ways, so we can take "2" red, "1" yellow and "3" white balls by "\binom{8}{2}\binom{3}{1}\binom{9}{3}" ways.

In general we can take "6" of "20" balls by "\binom{20}{6}" ways.

So our probability is "\frac{\binom{8}{2}\binom{3}{1}\binom{9}{3}}{\binom{20}{6}}=\frac{\frac{8!}{2!6!}\cdot\frac{3!}{2!1!}\cdot\frac{9!}{3!6!}}{\frac{20!}{6!14!}}=\frac{28\cdot 3\cdot 84}{38760}=\frac{294}{1615}"

Answer: "\frac{294}{1615}"